OpenStudy (anonymous):

x^2-9/x+2 over x-3/x-2 answer is (x-2) (x+3) over x+2 How do u solve? Step by step

OpenStudy (anonymous):

x^2-3=(x-3)(x+3) (x-3)(x+3)/(x+2)*(x-2)/(x-3) (x-3)(x+3)(x-2)/(x+2)(x-3) (x-3)cancel now we have: (x-2) (x+3) /( x+2)

OpenStudy (anonymous):

Sorry I was wrong number 9 x^2-9=(x-3)(x+3) (x-3)(x+3)/(x+2)*(x-2)/(x-3) (x-3)(x+3)(x-2)/(x+2)(x-3) (x-3)cancel now we have: (x-2) (x+3) /( x+2)

OpenStudy (anonymous):

It all makes sense up until the last part did u just flip the x-2 and the x+2?

OpenStudy (anonymous):

yes time and flip